OpenStudy (anonymous):

What is the value of x? 10 + 4 + 12x - 16

OpenStudy (anonymous):

@jennithemeani can u help me please

OpenStudy (babylove18):

is there supposed to be an = sign some were ?

OpenStudy (babylove18):

because you cant find x with out an = sign

OpenStudy (anonymous):

oops its its 10x + 4 = 12x - 16

OpenStudy (babylove18):

10x+4=12x-16 add 16 to both sides 10x+20=12x subtract 10x from both sides 20=2x divide both sides by 2 10=x

OpenStudy (anonymous):

thx

OpenStudy (tamara7301):

i agree with babylove18

OpenStudy (babylove18):

your welcome